### v1

# Token-Aware Coding Flow: A Study with Nano Surge in Reasoning Model thanks: * Corresponding authors

Abstract
With the widespread application of large-scale language models (LLMs) in software engineering, the Chain of Thought (CoT) approach has emerged as a crucial tool for driving automated code generation and optimization. However, despite the significant success of CoT methods in generating high-quality code, the issue of token inflation during the reasoning process remains a formidable challenge to model performance and efficiency, particularly when dealing with complex code smells. Code smells not only affect the maintainability and scalability of code but also significantly increase the computational burden during LLM inference, leading to excessive token consumption and, consequently, reduced reasoning efficiency. This paper introduces an innovative Token-Aware Coding Flow method, aimed at addressing the token inflation problem caused by smelly code in the CoT process. Through experimentation, we validate the synergistic effect of code refactoring and prompt engineering strategies, demonstrating that after eliminating code smells, token consumption during model inference is significantly reduced. The experimental results show that refactored code, while maintaining functional consistency, can reduce token consumption by up to 50%. Additionally, by explicitly prompting the type of code smells in the prompt and incorporating strategies such as context awareness and role constraints, we further optimize the reasoning process, achieving a 24.5% to 30% reduction in token consumption. These optimizations not only significantly enhance the model’s reasoning efficiency and improve code generation quality but also provide new insights for addressing performance bottlenecks in complex code generation tasks.
Index Terms:
code smell, CoT, token consumption
I Introduction
Figure 1: Evolution of software development paradigms. (1) Traditional coding paradigm: human programmers manually translate requirements into code. (2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort. (3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
With the rapid proliferation of large language models (LLMs), such as GPT-o3 and DeepSeek-R1[1], in the field of software development[2], traditional software development paradigms have undergone profound changes(Fig. 1). In the past, software functionality was implemented solely through manual programming; however, it is now gradually shifting toward human-machine collaboration[3], where code generation and optimization are assisted or directly achieved through the ”Chain of Thought” (CoT) approach[4]. This new development paradigm significantly improves both development efficiency and code quality[5]. However, existing research primarily focuses on LLM performance under ideal conditions, such as code that is well-structured and well-designed[6], neglecting the impact of ”code smells” that are commonly present in industrial settings due to legacy issues or insufficient experience.
Code smells[7] are pervasive in industrial software systems and typically manifest as complex, redundant logical structures, ambiguous naming conventions, excessively long functions or parameter lists, and overly coupled class structures. While these issues may not immediately cause software failures, they severely impair code readability and maintainability in the long run, and may even lead to latent defects[8]. Previous studies have shown that human developers face an additional cognitive burden when dealing with these smelly codes, leading to reduced development efficiency[9]. However, it remains unclear how such smelly code affects model inference efficiency and token consumption, especially in the emerging LLM-driven development paradigm.
Preliminary practical observations suggest that large language models exhibit noticeable ”token inflation”[10] and ”smell regurgitation”[11] phenomena when processing code snippets with smells(Fig. 2). Specifically, the model may repeatedly validate and reason about a particular code fragment or logical unit during inference, resulting in an overall redundant reasoning process that significantly increases token consumption[12]. This not only raises the economic cost of using large models [13] but also severely limits their applicability in resource-constrained environments (e.g., mobile platforms and real-time applications). Therefore, a deeper understanding of and strategies to reduce the negative impact of code smells on model inference efficiency is of great practical significance.
Figure 2: An example of redundant reasoning caused by confusion between C (row-major) and Fortran (column-major) memory layout. The user request involves reshaping a 4D CNN mask into a 2D time series mask compatible with RNNs, which requires correct alignment with Fortran-order expectations. However, the reasoning model repeatedly revisits the same subproblem—how elements are ordered in memory and how reshaping should be done—without converging on a solution. This looped verification, highlighted by repeated statements and backward arrows, illustrates the reasoning inefficiency when the model lacks concrete assumptions about memory layout or reshape semantics.
To further understand and attempt to address the above challenges, this paper proposes a method called Token-Aware Coding Flow. We first constructed a dataset containing ten common Java code smells, as well as a clean code dataset, based on a portion of the CodeXGLUE[14] Text-Code dataset, using the tree_sitter syntax tree construction tool[15]. This dataset was used to observe the token consumption differences when the DeepSeek-R1 model is faced with ”clean code” versus ”smelly code.” The results show that ”smelly code” consumes significantly more tokens per unit time than ”clean code,” highlighting the negative impact of code smells on model inference efficiency[16].
Building on this, we designed a series of exploratory experiments to systematically investigate the impact of different optimization strategies on token consumption. This study focuses on the impact of smelly code on inference models and introduces the ”Token-Aware Coding Flow” method. Through a series of carefully designed empirical experiments, we delve into the token consumption characteristics of smelly code during the Chain of Thought process and propose corresponding optimization strategies[17]. Specifically, this study aims to answer the following research questions (RQs):
  • RQ1: What are the specific impacts of smelly code on the reasoning process of large models (DeepSeek-R1) compared to clean code?
  • RQ2: How does code refactoring (eliminating code smells) impact token consumption during model inference?
  • RQ3: Are there significant differences in token consumption caused by different types of code smells?
  • RQ4: Does explicitly indicating the type of code smells in the model prompt help reduce token consumption during inference?
  • RQ5: In addition to code refactoring, what other effective prompt engineering strategies can further reduce token consumption when reasoning with smelly code?
Through experiments, we obtained the following key findings:
  • Finding1: Compared to clean code, the model consumes approximately 32% more tokens per unit time when processing smelly code.
  • Finding2: After eliminating code smells through refactoring, and ensuring functional consistency, token consumption during inference can be reduced by approximately 50%.
  • Finding3: The impact of different types of code smells varies significantly, with design and structural smells imposing a greater burden on token consumption than naming-related smells.
  • Finding4: Explicitly indicating code smell types in the prompt can reduce token consumption by approximately 24.5%.
  • Finding5: Combining context awareness, role constraints, and cost-sensitive prompt strategies can further reduce token consumption by 2% to 25%.
Based on the above research questions and findings, the main contributions of this paper are as follows:
  • Contribution1: We are the first to explore the impact of code smells on token consumption during model inference.
  • Contribution2: We validate the potential of code refactoring in reducing token consumption and provide a reusable method for automating the refactoring of smelly code.
  • Contribution3: We reveal the differential impact of various types of code smells on the model’s inference process.
  • Contribution4: We propose and validate effective prompt engineering strategies for smelly code handling.
II Context
II-A Code Smell
”Code Smell” refers to code characteristics that, while functional, may indicate underlying issues such as poor quality, maintenance difficulties, and scalability problems. These smells, like duplicated code, long methods, and excessive parameters[18], often result from bad programming practices and can negatively affect performance and readability. While they do not always cause immediate errors, they can complicate future maintenance. Developers often address code smells through refactoring, though manual inspection in large projects can be time-consuming and error-prone. Automated methods using machine learning and pre-trained models are increasingly being explored to improve the detection and elimination of code smells.
II-B Reasoning Module
The reasoning module is essential in intelligent code analysis and generation systems. It uses logical reasoning to interpret input (like code or task descriptions) and generate appropriate outputs. In code generation and repair tasks, it helps the model understand code structure, identify issues, and generate conforming code. With advancements in deep learning, reasoning modules have evolved from rule-based systems to those powered by large pre-trained models like GPT-3 and Codex. These models enhance reasoning accuracy and efficiency, making them crucial for tasks like automated code repair and generation(Fig. 3)[19].
Figure 3: Illustration of the LLM-driven code generation workflow in real-world development scenarios. The process begins with diverse input sources, such as natural language (NL), code translation tasks, bug reports, partial implementations, or test cases. The large language model (LLM) engages in multi-turn reasoning by interpreting the input, thinking step-by-step (e.g., via inner monologue), and generating candidate code. If the generated code fails to meet requirements, the model rethinks the problem by refining prompts and iterating on its reasoning. The process concludes with the production of executable code that satisfies the development goal.
III Experimental Insights
III-A Research Questions
To explore the impact of code smells on the reasoning process of large models, we designed a series of experiments aimed at addressing the following research questions:
  • RQ1: What are the specific impacts of smelly code on the reasoning process of large models (DeepSeek-R1) compared to clean code?
  • RQ2: How does code refactoring (eliminating code smells) impact token consumption during model inference?
  • RQ3: Are there significant differences in token consumption caused by different types of code smells?
  • RQ4: Does explicitly indicating the type of code smells in the model prompt help reduce token consumption during inference?
  • RQ5: In addition to code refactoring, what other effective prompt engineering strategies can further reduce token consumption when reasoning with smelly code?
Through these questions, we aim to gain a deeper understanding of the impact of smelly code on large model inference and explore the potential improvements in reasoning efficiency with different optimization strategies[20]. First, RQ1 primarily focuses on whether smelly code causes large models to consume more tokens during inference, which is crucial for evaluating the model’s ability to handle non-standard code. RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency. RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption. RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption. Finally, RQ5 explores whether other effective prompt engineering strategies, such as context awareness, role constraints, and cost-sensitive strategies, can further optimize reasoning efficiency beyond code refactoring. By addressing these questions, we aim to both validate the specific impact of code smells on large model reasoning performance and provide experimental evidence for improving inference efficiency through refactoring and optimized prompt design.
III-B Tasks
In this study, we designed a series of evaluation tasks to systematically explore the impact of code smells on the reasoning process of large models [21] and to test the effectiveness of different optimization strategies. The following are the specific tasks set for each research question.
III-B1 Observational Experiment
Datasets: We used the Java dataset (Text-To-Code) from CodeXGLUE[14] for the experimental data, which includes 300 samples of smelly code and 300 samples of clean code. These datasets were used to analyze the behavior of the DeepSeek-R1 reasoning model when processing these codes, particularly the differences in token consumption and time consumption.
Metrics: We recorded the total number of tokens consumed by the model for each code and the time taken for reasoning. The token consumption per unit time was calculated as the core metric to assess the efficiency of large models when handling different types of code. Additionally, the generated code was evaluated based on functionality correctness, readability, robustness, and other indicators.
III-B2 Refactoring[22] and Optimization Experiment
Datasets: In this experiment, we used the dataset based on smelly code and applied DeepSeek-R1 for code refactoring to generate refactored code (rf_code). The refactored code was used for chain processing to compare its reasoning efficiency with the original smelly code and clean code.
Metrics: We recorded the number of tokens consumed in each processing step and calculated the Halstead complexity and cyclomatic complexity. By assessing the token consumption per unit of Halstead complexity and per line of code, we evaluated the efficiency improvements of the refactored code during model inference. Further, the accuracy of the generated code was assessed through functionality consistency scores (e.g., CodeBLEU) to ensure that the refactored code maintains functional consistency with the original code.
III-B3 Code Smell Type Comparison Experiment
Datasets: We selected code samples from 10 different types of code smells in the smelly code dataset, with 50 samples from each type, totaling 500 samples. These samples were used to analyze the impact of different types of code smells on token consumption during model inference.
Metrics: By comparing the token consumption during reasoning for different types of code smells, we calculated the token consumption per unit time for each smell type. The differences in the impacts of design, structural, and naming-related smells were further analyzed. All generated code was evaluated for functional consistency to ensure that the model-generated code met the expected functionality during reasoning.
III-B4 Explicit Smell Type Prompt Experiment
Datasets: This experiment was conducted on the previously used smelly code dataset, primarily comparing the token consumption differences when explicitly indicating code smell types in the prompt versus not indicating them.
Metrics: We recorded the number of tokens consumed in each experimental step and calculated token consumption per unit time. After generating the code, functionality consistency was assessed using metrics like CodeBLEU. Additionally, the effect of explicitly indicating code smell types was evaluated based on the model’s reasoning time, token consumption, and output quality.
III-B5 Prompt Optimization Strategy Experiment
Datasets: In this experiment, we tested different prompt optimization strategies on the smelly code dataset[23], analyzing the effects of Context Awareness, Responsibility Tuning, and Cost-Sensitive strategies.
Metrics: We recorded token consumption during the reasoning process for each strategy and calculated the Halstead complexity and cyclomatic complexity for each code. By comparing token consumption per unit complexity and per line of code, we assessed the impact of the three prompt optimization strategies on reasoning efficiency. All experiments were also evaluated for functionality consistency and code generation quality to ensure the generated code was accurate[16].
III-C Implementation
This study’s experimental implementation is built upon existing LLMs and open‑source code‑refactoring tools, with all tasks orchestrated via API calls to guarantee reproducibility and efficiency[24]. Specifically, we employ DeepSeek‑R1 as the inference engine and integrate Tree‑Sitter for code‑smell detection and automated refactoring. The implementation proceeds as follows(Fig. 4):
Figure 4: Experiment Design Process. The diagram shows the steps involved in extracting requirements, generating code, refactoring code, and comparing semantic and code similarities. The flow demonstrates the interaction between different components, including code refactoring and requirement extraction, with the evaluation steps represented through similarity comparison.
  • Data Processing [20] and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed with Tree‑Sitter to detect and classify instances of code smell. Leveraging Tree‑Sitter’s high‑performance syntax analysis, we annotated ten common smell categories, thereby constructing parallel “smelly” and “clean” code datasets. For each smelly sample, we invoked the DeepSeek‑R1 API to execute the reasoning task, recording both token consumption and inference latency. We then applied DeepSeek‑R1’s refactoring API to eliminate smells—producing rf_code—and compared the inference efficiency (tokens consumed and time taken) between original and refactored versions.
  • Prompt Engineering and Optimization[25]:To further reduce token overhead, we evaluated three prompt‑level strategies: context awareness, role constraints, and cost sensitivity. Each strategy was realized by modifying the model’s input prompt, thereby influencing the reasoning trajectory and token usage. For every experimental condition, we issued the corresponding API calls to adjust generation parameters and logged token counts, inference durations, and the resulting code quality.
  • Automation, Logging, and Analysis:All API interactions were fully automated via Python scripts, ensuring consistent and efficient execution. Each experiment’s outputs—including raw logs of token usage, timing, and generated code—were aggregated and processed by custom analysis scripts. These scripts perform data cleansing, statistical aggregation, and visualization, enabling systematic comparison across differing experimental configurations.
IV Experimental Results
IV-1 RQ1: Smelly Code vs. Clean Code
Experimental Design.
We randomly sampled 30 instances from each of the ten individual code‑smell datasets and aggregated them into a single file, smelly_code.jsonl. We then randomly selected 300 examples from the clean‑code dataset, producing clean_code.jsonl. Because the two collections differ substantially in content and structure, we employ the Time‑Scaled Token Consumption metric to evaluate token expenditure during inference. Using the DeepSeek‑R1 model, each sample is assessed across six dimensions—functional correctness, readability, robustness, maintainability, extensibility, and security—while we record the inference procedure, evaluation outcomes, and end‑to‑end execution time. From these logs, we compute the total tokens consumed during both the reasoning and output‑generation phases.
Analysis.
In our analysis, we first compare token consumption per unit time between clean code and smelly code. Directly contrasting absolute token counts would misrepresent efficiency, since the code samples vary greatly in complexity and structure. By normalizing to tokens per unit time, we obtain a more precise measure of model throughput across code types. Statistical evaluation reveals that smelly code consistently demands higher token expenditure(Fig. 5)(Tab. I). This elevated consumption stems from the increased complexity of smelly code: the model must undertake additional reasoning steps and verifications to resolve intricate structural and logical relationships. Conversely, clean code—with its clearer organization—allows the model to complete inference with fewer redundant operations.
Figure 5: Comparison of time-scaled token consumption between clean code and smelly code.
TABLE I: Descriptive Statistics of Time-Scaled Token Consumption
Statistic Clean Code Smelly Code
Count 300 300
Mean 24.44 33.20
Std. Dev. 12.35 18.63
Min 0.61 0.83
25% Quartile 18.55 25.30
Median (50%) 28.13 33.30
75% Quartile 31.68 41.90
Max 58.97 120.16
Findings – Smelly Code vs Clean Code The results show that smelly code leads to significantly higher token consumption during inference than clean code. This is mainly attributed to structural complexity such as redundant logic and poor naming, which forces the model to perform repeated validations. These findings suggest that code smells not only harm maintainability but also increase the inference cost of large models, highlighting the importance of code quality for efficient model reasoning.
IV-A RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
Experimental Setup.
To answer RQ2, we designed an experiment to investigate how code refactoring—specifically the elimination of code smells—affects token consumption during model inference. We first used the DeepSeek‑R1 inference model on both smelly_code.jsonl and clean_code.jsonl to establish baseline token‑usage data. Next, we applied automated refactoring to the smelly code samples, removing common smells (e.g., long methods, duplicated logic) to generate a refactored code set (rf_code). We then re‑ran inference on these refactored samples and compared their token consumption to the original smelly code. To ensure functional consistency, we employed CodeBLEU and other metrics to verify that the refactored code preserved the original behavior.
Analysis.
During analysis, we recorded both the inference duration and token consumption for each trial and computed token consumption per unit time. Under the functional consistency constraint, refactored code is expected to be leaner, reducing unnecessary complexity and redundant patterns. Our results show that after refactoring, the DeepSeek‑R1 model performs inference more efficiently: despite maintaining equivalent functionality(Tab. III), the streamlined code requires significantly fewer tokens(Fig. 6)(Tab. II). In particular, refactoring complex smells—such as excessively long functions and duplicated segments—substantially reduced the number of reasoning steps and validations the model must perform.
Figure 6: Comparison of complexity-normalized token consumption before and after code refactoring.
TABLE II: Descriptive Statistics of Token per Complexity
Statistic Original Code Refactored Code
Count 300 300
Mean 0.1015 0.0576
Std. Dev. 0.1732 0.0816
Min 0.0020 0.0020
25% Quartile 0.0144 0.0094
Median (50%) 0.0334 0.0239
75% Quartile 0.1180 0.0723
Max 1.5613 0.5836
TABLE III: Summary of CodeBLEU and Docstring Similarity Statistics (Mean Values Only)
Category Metric Mean
CodeBLEU code & rf code 0.5224
rf code & rf gc code 0.1803
Docstring code & rf code 0.8604
code & rf gc code 0.7356
rf code & rf gc code 0.7364
Findings – Refactoring Impact Code refactoring yields a substantial reduction in token consumption during inference while preserving functional correctness. Specifically, the refactored code consumes approximately 50% fewer tokens compared to the original smelly code. By removing redundant structures and simplifying complex fragments, refactoring not only enhances code quality but also markedly improves inference efficiency. These results suggest that systematic elimination of code smells is a critical strategy for reducing computational overhead in large‑model reasoning.
IV-B RQ3: Impact of Different Types of Code Smells on Token Consumption
Experimental Design / Setup
To answer RQ3, we designed an experiment to investigate how different types of code smells affect token consumption in the DeepSeek‑R1 inference model. We extracted from smelly_code.jsonl ten common smell categories—complicated_regex_expression, parameter_list_too_long, binary_operator_in_name, complicated_boolean_expression, etc.—with 30 code snippets per category (300 total). We ran DeepSeek‑R1 inference separately on each group, recording the tokens consumed for each snippet. By computing the tokens consumed per unit time for each smell type, we identified which categories impose the greatest inference burden.
Analysis
The first graph(Fig. 7) and the table(Tab. IV))present the average growth rate of token consumption across various code smell types. It is evident that most of the code smells, such as binary_operator_in_name, complicated_boolean_expression, and complicated_regex_expression, lead to a significant increase in token consumption. These types of smells are often characterized by complex logical constructs that demand deeper reasoning from the model, resulting in more tokens being processed[26]. Conversely, the smells func_name and too_long_parameter_list have a relatively lower impact on token growth. This suggests that these types of smells are simpler and less likely to introduce complexity in reasoning, causing fewer tokens to be consumed.
The second graph(Fig. 8) categorizes code smells into four groups: Naming, Expression, Structure, and Design. The Expression category stands out with the highest average token consumption growth, followed by Structure. These categories typically involve complex logical constructs or convoluted code structures, which require more tokens for the model to parse and reason about. In contrast, the Naming and Design categories exhibit lower token consumption growth. Issues like poor naming conventions and design flaws might affect code readability and maintainability, but they do not significantly complicate the reasoning process, leading to lower token usage by the model.
Figure 7: Average token growth rate per code smell type.
Figure 8: Average token growth rate by code smell category.
TABLE IV: Average Token Growth Rate by Smell Type
Smell Type Avg. Growth Rate
binary_operator_in_name 0.3424
complicated_boolean_expression 0.3361
complicated_regex_expression 0.3221
cyclomatic_complexity 0.3285
func_name 0.1592
loops 0.3111
mutation_too_much 0.3321
too_long_parameter_list 0.1942
primitive_obsession 0.2963
too_long 0.3089
TABLE V: Comparison of Code Equivalence Before and After Refactoring (Mean Values Only)
Category Metric Mean
Source Code code & gc code 0.1659
Refactored Code code & rf code 0.5208
rf code & rf gc code 0.1773
Source Code Similarity code & gc code 0.7470
Refactored Code Similarity code & rf code 0.8610
code & rf gc code 0.7295
rf code & rf gc code 0.7340
Findings – Smell Type Impact Our study, conducted with a 70% code functionality similarity threshold(Tab. V), reveals significant variance in token consumption across different code smell categories. Complex code smells—namely complicated_regex_expression, parameter_list_too_long, and complicated_boolean_expression—incur the highest token costs due to increased code complexity and the need for additional reasoning steps. Simpler smells, such as binary_operator_in_name and func_name, result in comparatively minor token overhead. These findings suggest that prioritizing the removal of design and structural smells can maximize inference efficiency and substantially reduce computational overhead.
IV-C RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[27]
Experimental Design / Setup
To address RQ4, we designed an experiment to investigate whether explicitly annotating code smell types in the model prompt effectively reduces token consumption during inference. We compared two conditions on the same smelly_code.jsonl dataset: (1) No annotation, using the raw code as input; and (2) Explicit annotation, where the prompt was augmented with statements such as “the code contains complex regular expressions” or “the code has an excessively long parameter list.” For each condition, we invoked DeepSeek‑R1 to perform inference and recorded both total token usage and inference time. We then computed token consumption per unit time to facilitate a direct comparison between annotated and unannotated prompts.
Analysis
Our study, conducted with a 70% code functionality similarity threshold (Tab. VII). During analysis, we examined the token consumption profiles under each prompting condition. Results (Fig. 9)(Tab. VI) indicate that adding explicit smell annotations to the prompt yields a more efficient reasoning process, with a marked reduction in token usage. Explicit annotations supply the model with rich contextual cues about potential issues, enabling it to bypass redundant reasoning and validation steps. In contrast, without smell hints, the model must engage in multiple verification cycles to detect and interpret hidden code complexities, thereby inflating token consumption.
Figure 9: Comparison of total token consumption for smelly code with and without explicit prompt tips
TABLE VI: Descriptive Statistics of Total Token Consumption for Smelly Code (With vs. Without Prompt Tips)
Statistic No Tips With Tips
Count 300 300
Mean 5876.49 4431.19
Std. Dev. 3172.77 2578.35
Min 1095.00 981.00
25% Quartile 3419.50 2312.50
Median (50%) 5351.00 3685.00
75% Quartile 7985.75 5820.00
Max 16440.00 15045.00
TABLE VII: Comparison of CodeBLEU and Similarity Metrics with and without Tips
Metric Mean (With Tips) Mean (No Tips)
CodeBLEU 0.1585 0.1665
Code Similarity 0.7248 0.7619
Findings – Explicit Smell Prompt Explicitly indicating code smell types in the model prompt reduces token consumption by approximately 24.5% compared to unannotated prompts. The provision of clear smell-specific cues allows the model to focus its reasoning on relevant code fragments, eliminating unnecessary validation steps and lowering computational overhead. This strategy thus significantly enhances inference efficiency and mitigates resource usage in large‑model code reasoning.
IV-D RQ5: Prompt Engineering Strategies beyond Refactoring
TABLE VIII: Mean Values of Key Metrics under Different Strategies
Strategy Comp‑Norm Code Line‑Scaled Code Comp‑Norm Token Line‑Scaled Token
Base 0.1015 46.7589 0.6026 277.7973
Context 0.0858 39.3372 0.4979 229.0921
Func 0.0891 40.9848 0.5357 241.2648
Total 0.0981 44.9317 0.4760 220.8598
DevOps 0.0852 38.5346 0.4832 218.3642
QAer 0.0868 40.4725 0.5362 238.8191
SEer 0.0771 35.5723 0.5535 241.9891
AbsCost 0.0815 37.1547 0.4746 202.5059
RelCost 0.0859 38.7631 0.4038 178.0056
Comb1 0.1001 44.5918 0.5705 245.4126
Comb2 0.0981 44.1189 0.5301 227.1736
TABLE IX: CodeBLEU and Docstring Similarity Statistics
Category Method Mean CodeBLEU Mean Code-and-GC-Code Similarity
Baseline - 0.1803 0.7364
Context Awareness Code Context 0.1783 0.7428
Code Function 0.2283 0.7668
All Information 0.2464 0.7828
Responsibility Tuning SEer 0.1593 0.7210
QAer 0.1590 0.7369
DevOps 0.1570 0.7275
Cost Sensitive Absolutely 0.1605 0.7011
Relatively 0.1799 0.6988
Combination All Methods 0.2513 0.7635
All Methods 0.2505 0.7665
Experimental Design / Setup
To address RQ5, we designed a series of experiments(Fig. 10) aimed at evaluating three prompt engineering strategies—Context Awareness, Responsibility Tuning, and Cost Sensitivity—and their effect on reducing token consumption during the model’s reasoning over smelly code. Each strategy modifies the input prompt to guide the model’s cognitive process, driving a more efficient and focused inference, while minimizing unnecessary token expenditure. Specifically:
  • Context Awareness: By infusing the prompts with contextual details such as file path, module or function names, and surrounding code snippets, we enhanced the model’s comprehension of the code’s role within the broader system. This enriched context directs the model’s focus to key segments of the code, thereby avoiding redundant inference over irrelevant portions.
  • Responsibility Tuning: Through the strategic assignment of roles—such as “software engineer,” “QA engineer,” or “DevOps engineer”—we framed the model’s reasoning process to align with a specific perspective. This role-induced focus curtailed extraneous reasoning steps and unnecessary pathways, leading to a more efficient processing of the code.
  • Cost Sensitivity: We introduced explicit limits on the length of the generated output or the token count (e.g., maximum token budget). By constraining the token size, we controlled the model’s computational cost, particularly when generating larger code segments, effectively reducing token expenditure without sacrificing the integrity of the generation.
All experiments were performed on the same dataset of 300 smelly code snippets (smelly_code.jsonl). For each condition, we invoked DeepSeek‑R1, recording the total token consumption and calculating token consumption per unit time. Results were compared against a baseline prompt, which did not include any specific optimization strategies.
Analysis
We analyzed token consumption across different prompt conditions by calculating tokens per unit time for each strategy. Our key observations include(Tab. VIII):
  • Context Awareness: By enriching the prompts with contextual information, the model was able to bypass irrelevant logic and focus on the core components of the code. On average, token consumption decreased by 15–20%, with more significant reductions observed in longer, more complex code fragments.
  • Responsibility Tuning: Assigning roles to the model significantly honed its focus on task‑relevant reasoning steps. This strategy yielded a 10–15% reduction in token consumption, particularly in large‑scale codebases, where the model’s focus was sharpened by the specified role.
  • Cost Sensitivity: Imposing limits on output length or token budget led to a 20–30% reduction in token expenditure. However, overly stringent constraints sometimes led to the truncation of essential code elements, resulting in reduced functional correctness and lower accuracy in the output.
Figure 10: Nano Surge on Token Consumption
Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning emerged as the most effective strategies for reducing token consumption beyond refactoring(Tab. IX). Context Awareness decreased token usage by 15–20%, especially on lengthy or complex code. Responsibility Tuning achieved a 10–15% reduction by framing the model’s reasoning role. Cost Sensitivity, while reducing tokens by 20–30%, risked incomplete code generation under strict limits. These results suggest that combining context-rich prompts with precise role framing offers substantial efficiency gains without compromising functionality. Furthermore, the optimizations applied through these strategies also led to improvements in code quality metrics. Specifically, the codeBLEU score increased by 38%, and function similarity was enhanced by 4%, compared to the baseline performance. These improvements indicate that not only were token consumption and reasoning efficiency optimized, but the overall quality of the generated code also saw significant gains, further validating the efficacy of these prompt engineering strategies.
V Related Work
Over recent years, with the widespread application of large language models (LLMs) in natural language processing (NLP)[3], automated code generation and optimization have also made significant strides[5]. Research in this area can be broadly categorized into three directions:
V-A Code Smell Detection and Refactoring
Code smells are syntactic or structural code features that do not immediately impair program functionality but degrade readability, maintainability, and extensibility[7][8][9]. Tufano et al. (2017)[28] proposed an automated smell detection approach by analyzing code structures and patterns to identify potential smell instances. Lanza (2003)[29] further investigated the impact of code smells on program maintainability and introduced corresponding refactoring strategies. Many refactoring tools combine rule‑based and machine learning techniques to automatically detect and repair smells at scale[7]. Nevertheless, efficiently handling complex smells remains challenging, particularly in deep learning–based inference where token consumption and computational overhead can escalate.
V-B LLM‑based Code Generation and Optimization
The advent of pre‑trained language models has spurred extensive research into LLM‑driven code generation. Models such as OpenAI’s GPT series and Codex, pre‑trained on large code corpora, demonstrate strong capabilities in code completion, synthesis, and repair. The Transformer architecture of Vaswani et al. (2017) laid the foundation for these tasks, and subsequent variants such as BERT and GPT have further advanced code generation quality. While LLMs capture syntactic and semantic patterns effectively, their inference often suffers from token inflation, especially when processing complex or smelly code. To mitigate this, researchers have explored integrating automated refactoring with LLM training to jointly optimize code quality and inference efficiency.
V-C Integration of Code Quality and Inference Efficiency
Recently, there has been growing interest in coupling code quality improvements with inference efficiency in pre‑trained models. Pre‑trained models like DeepCode and CodeBERT exhibit strong performance in code quality assessment and automated repair. Empirical studies indicate that higher code quality not only enhances maintainability but also reduces inference costs. For example, Zhang et al. (2020) proposed a deep learning–based optimization method that produces functionally equivalent code while significantly lowering token consumption during inference. Building upon this work, our study investigates how code smell removal and prompt engineering jointly reduce token overhead, particularly on complex code fragments.
VI Conclusion and Future Work
This study investigated strategies to optimize the reasoning pipeline of large language models (LLMs) in order to reduce token consumption when processing code with smells[16][14]. We proposed a combined approach of automated code refactoring and prompt engineering[22], and empirically demonstrated that these strategies significantly enhance inference efficiency[13]. Our experimental results show that code smells substantially inflate token usage during inference, whereas eliminating smells via refactoring and refining prompts markedly reduces computational overhead[8][9]. Specifically, code refactoring alone decreased token consumption by approximately 50%, explicit smell annotations in the prompt yielded a 24.5% reduction, and prompt engineering techniques such as Context Awareness and Responsibility Tuning provided further savings.
Despite these advances, several limitations and avenues for future work remain. First, our experiments focused exclusively on Java and the DeepSeek‑R1 model; future research should extend our methodology to other programming languages and inference engines to assess its generalizability. Second, while refactoring and prompt optimizations proved effective, complex code‐smell scenarios may still challenge inference efficiency[13]. Enhancing model reasoning capabilities for diverse programming tasks and more intricate smell types warrants deeper investigation.
Moreover, current work has largely concentrated on token consumption and inference speed[19]. Future studies should explore the trade‐off between generation quality and efficiency—for example, how to maintain code correctness, maintainability, and extensibility while further reducing token usage[27]. Integrating automated refactoring with deep learning–based optimization could also offer practical guidance for improving code quality in large codebases and advancing intelligent programming tools.
In summary, this research provides new insights into token optimization in code generation and reasoning, highlighting the pivotal role of smell elimination and prompt refinement. As intelligent coding assistants evolve, we anticipate that more granular reasoning optimizations and strategic prompt designs will enable next‐generation code generation systems to meet real‐world demands with greater efficiency and precision.

### v2

@@
-Token-Aware Coding Flow: A Study with Nano Surge in Reasoning Model thanks: * Corresponding authors
+Optimizing Token Consumption in LLMs: A Nano Surge Approach for Code Reasoning Efficiency thanks: * Corresponding authors
 Abstract
-With the widespread application of large-scale language models (LLMs) in software engineering, the Chain of Thought (CoT) approach has emerged as a crucial tool for driving automated code generation and optimization.
-However, despite the significant success of CoT methods in generating high-quality code, the issue of token inflation during the reasoning process remains a formidable challenge to model performance and efficiency, particularly when dealing with complex code smells.
-Code smells not only affect the maintainability and scalability of code but also significantly increase the computational burden during LLM inference, leading to excessive token consumption and, consequently, reduced reasoning efficiency.
-This paper introduces an innovative Token-Aware Coding Flow method, aimed at addressing the token inflation problem caused by smelly code in the CoT process.
-Through experimentation, we validate the synergistic effect of code refactoring and prompt engineering strategies, demonstrating that after eliminating code smells, token consumption during model inference is significantly reduced.
-The experimental results show that refactored code, while maintaining functional consistency, can reduce token consumption by up to 50%.
-Additionally, by explicitly prompting the type of code smells in the prompt and incorporating strategies such as context awareness and role constraints, we further optimize the reasoning process, achieving a 24.5% to 30% reduction in token consumption.
-These optimizations not only significantly enhance the model’s reasoning efficiency and improve code generation quality but also provide new insights for addressing performance bottlenecks in complex code generation tasks.
+With the increasing adoption of large language models (LLMs) in software engineering, the Chain of Thought (CoT) reasoning paradigm has become an essential approach for automated code repair.
+However, the explicit multi-step reasoning in CoT leads to substantial increases in token consumption, reducing inference efficiency and raising computational costs, especially for complex code repair tasks.
+Most prior research has focused on improving the correctness of code repair while largely overlooking the resource efficiency of the reasoning process itself.
+To address this challenge, this paper proposes three targeted optimization strategies: Context Awareness, Responsibility Tuning, and Cost Sensitive.
+Context Awareness guides the model to focus on key contextual information, Responsibility Tuning refines the structure of the reasoning process through clearer role and responsibility assignment, and Cost Sensitive incorporates resource-awareness to suppress unnecessary token generation during inference.
+Experiments across diverse code repair scenarios demonstrate that these methods can significantly reduce token consumption in CoT-based reasoning without compromising repair quality.
+This work provides novel insights and methodological guidance for enhancing the efficiency of LLM-driven code repair tasks in software engineering.
 Index Terms:
@@
 I Introduction
-Figure 1: Evolution of software development paradigms.
-(1) Traditional coding paradigm: human programmers manually translate requirements into code.
-(2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort.
-(3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
-With the rapid proliferation of large language models (LLMs), such as GPT-o3 and DeepSeek-R1[1], in the field of software development[2], traditional software development paradigms have undergone profound changes(Fig.
-1).
-In the past, software functionality was implemented solely through manual programming; however, it is now gradually shifting toward human-machine collaboration[3], where code generation and optimization are assisted or directly achieved through the ”Chain of Thought” (CoT) approach[4].
-This new development paradigm significantly improves both development efficiency and code quality[5].
-However, existing research primarily focuses on LLM performance under ideal conditions, such as code that is well-structured and well-designed[6], neglecting the impact of ”code smells” that are commonly present in industrial settings due to legacy issues or insufficient experience.
-Code smells[7] are pervasive in industrial software systems and typically manifest as complex, redundant logical structures, ambiguous naming conventions, excessively long functions or parameter lists, and overly coupled class structures.
-While these issues may not immediately cause software failures, they severely impair code readability and maintainability in the long run, and may even lead to latent defects[8].
-Previous studies have shown that human developers face an additional cognitive burden when dealing with these smelly codes, leading to reduced development efficiency[9].
-However, it remains unclear how such smelly code affects model inference efficiency and token consumption, especially in the emerging LLM-driven development paradigm.
-Preliminary practical observations suggest that large language models exhibit noticeable ”token inflation”[10] and ”smell regurgitation”[11] phenomena when processing code snippets with smells(Fig.
-2).
-Specifically, the model may repeatedly validate and reason about a particular code fragment or logical unit during inference, resulting in an overall redundant reasoning process that significantly increases token consumption[12].
-This not only raises the economic cost of using large models [13] but also severely limits their applicability in resource-constrained environments (e.g., mobile platforms and real-time applications).
-Therefore, a deeper understanding of and strategies to reduce the negative impact of code smells on model inference efficiency is of great practical significance.
 Figure 1: An example of redundant reasoning caused by confusion between C (row-major) and Fortran (column-major) memory layout.
@@
 This looped verification, highlighted by repeated statements and backward arrows, illustrates the reasoning inefficiency when the model lacks concrete assumptions about memory layout or reshape semantics.
-To further understand and attempt to address the above challenges, this paper proposes a method called Token-Aware Coding Flow.
-We first constructed a dataset containing ten common Java code smells, as well as a clean code dataset, based on a portion of the CodeXGLUE[14] Text-Code dataset, using the tree_sitter syntax tree construction tool[15].
-This dataset was used to observe the token consumption differences when the DeepSeek-R1 model is faced with ”clean code” versus ”smelly code.” The results show that ”smelly code” consumes significantly more tokens per unit time than ”clean code,” highlighting the negative impact of code smells on model inference efficiency[16].
-Building on this, we designed a series of exploratory experiments to systematically investigate the impact of different optimization strategies on token consumption.
-This study focuses on the impact of smelly code on inference models and introduces the ”Token-Aware Coding Flow” method.
-Through a series of carefully designed empirical experiments, we delve into the token consumption characteristics of smelly code during the Chain of Thought process and propose corresponding optimization strategies[17].
-Specifically, this study aims to answer the following research questions (RQs):
-  • RQ1: What are the specific impacts of smelly code on the reasoning process of large models (DeepSeek-R1) compared to clean code?
-  • RQ2: How does code refactoring (eliminating code smells) impact token consumption during model inference?
-  • RQ3: Are there significant differences in token consumption caused by different types of code smells?
-  • RQ4: Does explicitly indicating the type of code smells in the model prompt help reduce token consumption during inference?
-  • RQ5: In addition to code refactoring, what other effective prompt engineering strategies can further reduce token consumption when reasoning with smelly code?
-Through experiments, we obtained the following key findings:
-  • Finding1: Compared to clean code, the model consumes approximately 32% more tokens per unit time when processing smelly code.
-  • Finding2: After eliminating code smells through refactoring, and ensuring functional consistency, token consumption during inference can be reduced by approximately 50%.
-  • Finding3: The impact of different types of code smells varies significantly, with design and structural smells imposing a greater burden on token consumption than naming-related smells.
-  • Finding4: Explicitly indicating code smell types in the prompt can reduce token consumption by approximately 24.5%.
-  • Finding5: Combining context awareness, role constraints, and cost-sensitive prompt strategies can further reduce token consumption by 2% to 25%.
-Based on the above research questions and findings, the main contributions of this paper are as follows:
-  • Contribution1: We are the first to explore the impact of code smells on token consumption during model inference.
-  • Contribution2: We validate the potential of code refactoring in reducing token consumption and provide a reusable method for automating the refactoring of smelly code.
-  • Contribution3: We reveal the differential impact of various types of code smells on the model’s inference process.
-  • Contribution4: We propose and validate effective prompt engineering strategies for smelly code handling.
-II Context
-II-A Code Smell
-”Code Smell” refers to code characteristics that, while functional, may indicate underlying issues such as poor quality, maintenance difficulties, and scalability problems.
-These smells, like duplicated code, long methods, and excessive parameters[18], often result from bad programming practices and can negatively affect performance and readability.
-While they do not always cause immediate errors, they can complicate future maintenance.
-Developers often address code smells through refactoring, though manual inspection in large projects can be time-consuming and error-prone.
-Automated methods using machine learning and pre-trained models are increasingly being explored to improve the detection and elimination of code smells.
-II-B Reasoning Module
-The reasoning module is essential in intelligent code analysis and generation systems.
-It uses logical reasoning to interpret input (like code or task descriptions) and generate appropriate outputs.
-In code generation and repair tasks, it helps the model understand code structure, identify issues, and generate conforming code.
-With advancements in deep learning, reasoning modules have evolved from rule-based systems to those powered by large pre-trained models like GPT-3 and Codex.
-These models enhance reasoning accuracy and efficiency, making them crucial for tasks like automated code repair and generation(Fig.
-3)[19].
+The rapid proliferation of Large Language Models (LLMs) is driving a profound transformation within the software engineering domain.
+From code generation and documentation to intricate system design and defect analysis, LLMs exhibit unprecedented potential, fundamentally altering traditional software development practices.
+Among these advancements, Chain of Thought (CoT)[1] has garnered widespread attention as a strategy designed to enhance the reasoning capabilities of these models.
+By prompting LLMs to emulate a human-like, step-by-step problem-solving process, CoT significantly improves model accuracy and output quality when tackling complex software engineering tasks, such as generating intricate code, remediating deep-seated errors, and executing multi-step optimizations[2][3].
+This progress signifies a crucial shift for LLMs, from mere ”pattern matchers” to more sophisticated ”intelligent reasoners,” thereby unlocking new vistas for automated and intelligent software development[4].
+However, the enhanced capabilities afforded by CoT are accompanied by a significant and often overlooked challenge: a substantial increase in token consumption[5].
+The CoT mechanism inherently relies on the generation of detailed intermediate reasoning steps, which implies that both the length of input prompts and the volume of model-generated content can far exceed those of traditional, more succinct question-answering paradigms[6].
+This phenomenon, termed ”token inflation,”[7] has emerged as a critical bottleneck impeding the broader application of LLMs in the software engineering field.
+When LLMs are tasked with analyzing large-scale, complex modern software systems, or engaged in programming tasks requiring multiple iterative refinements, the token inflation precipitated by CoT becomes particularly acute[8].
+This not only imposes economic pressure on the cost-effectiveness of model utilization but also adversely impacts practical application efficiency.
+In real-world software development applications, cost control and operational efficiency remain paramount considerations[9].
+For all software development teams—ranging from large technology corporations and agile small-to-medium-sized enterprises to budget-constrained startups—the adoption of new technologies transcends mere enhancement of technical capabilities, necessitating a thorough evaluation of cost-effectiveness[6].
+Currently, a majority of LLM services are accessed via API calls, with charges typically levied based on the volume of tokens processed.
+While CoT can yield higher-quality outputs, the concomitant substantial increase in token consumption may render its use prohibitively expensive[8].
+Within modern software development pipelines, particularly in contexts such as mobile application development, embedded systems, and other resource-constrained environments, excessive token consumption can lead to considerable economic burdens, compelling many teams to exercise caution or curtail their reliance on CoT methodologies[1].
+Beyond direct economic costs, excessive token consumption engenders other detrimental effects, including prolonged model response times and heightened demands on computational resources.
+Such issues can impair the agility of software development processes and the immediacy of user experiences, especially in scenarios mandating rapid responses and efficient delivery[6].
+Consequently, achieving an optimal balance between harnessing the reasoning advantages of CoT and ensuring the economic viability and operational efficiency of LLM applications has surfaced as a significant challenge confronting the software engineering field[10].
+This study does not aim to propose a systemic solution; rather, it represents a preliminary exploration of this issue.
+Through the observation and analysis of token consumption phenomena within CoT processes, we aspire to furnish insights for future research, investigating methodologies to effectively curtail token consumption without a significant detriment to reasoning performance.
+This research endeavors to pave the way for the expanded application of LLMs in software engineering, particularly within contexts that prioritize developmental efficiency and resource optimization.
+Driven by widespread concerns regarding the token consumption induced by Chain of Thought (CoT) and its potential ramifications, this study undertakes an exploratory investigation.
+Our objective is to meticulously observe and comprehend the token dynamics of CoT when processing code of varying complexities within specific experimental settings[11].
+Furthermore, we aim to identify external intervention strategies capable of effectively mitigating such consumption.
+To ground our research in realistic and common software development contexts—scenarios where the intricacies of CoT’s reasoning pathways and token expenditure may be amplified—we concentrate our observational lens on how Large Language Models (LLMs) handle code exhibiting different intrinsic characteristics.
+Specifically, the presence and nature of ”code smells[12][13]” are employed as a tangible proxy and experimental variable for examining these phenomena.
+It is crucial to emphasize that our intent is not to present ”code smell elimination” as an end in itself, but rather to utilize it as an effective ”lens” through which to gain deeper insights into the broader challenge of CoT token optimization[14].
+Guided by this philosophy, we propose a conceptual methodological framework termed ”Token-Aware Coding Flow.” Based on this framework, we have conducted a series of preliminary experimental inquiries.
+These inquiries endeavor to offer initial insights into several key questions:
+  • What is the specific impact of intrinsic code characteristics (e.g., as studied through a comparative analysis of ”clean code” versus ”smelly code”) on token consumption during CoT reasoning processes?
+  • To what extent does pre-processing code (for instance, by refactoring to eliminate code smells, thereby simplifying the input furnished to CoT) demonstrate potential for reducing token consumption?
+  • Do different types of inherent code complexities (exemplified by various categories of code smells) lead to differentiated patterns in token consumption?
+  • When interacting with LLMs, does explicitly indicating potential code characteristics (such as the type of code smells) within the prompt help guide the CoT process toward more economical token usage?
+Our preliminary experiments reveal a clear correlation between the intrinsic characteristics of code and token consumption during Chain of Thought (CoT) reasoning.
+Compared to ”clean” code, code with ”smells” significantly increases token usage, while refactoring to remove such code smells can substantially reduce the tokens required for subsequent reasoning.
+Furthermore, different types of code complexity (such as various code smells) affect token consumption differently, with deeper logical or structural issues typically imposing a higher token burden.
+Explicitly indicating code characteristics in the prompt, combined with prompt engineering strategies, also demonstrates potential for optimizing token efficiency.
+Although this research is still in an exploratory phase and its conclusions require further validation, our initial results provide new perspectives on understanding and optimizing token consumption in CoT scenarios.Our main contributions of this work are as follows:
+  • For the first time, we systematically focus on resource (token) consumption in code generation and processing tasks, drawing attention to the economic and efficiency impact of CoT reasoning in LLM-based software engineering.
+  • We propose a novel approach to reduce resource consumption from outside the model, by leveraging code refactoring and prompt engineering techniques.
+This strategy offers practical and effective ways to optimize token usage without modifying model internals.
+  • We empirically validate that code smells lead to significantly higher token consumption, and demonstrate that external optimization strategies (such as code refactoring and prompt design) can effectively reduce this overhead.
+  • Our work provides empirical evidence and methodological inspiration for developing more cost-effective and efficient LLM-assisted software engineering workflows, contributing to the advancement of resource-aware intelligent development practices.
+II Introduce the New Coding Flow
+Figure 2: The Ladder of Programming and AI.
+(1) Traditional coding paradigm: human programmers manually translate requirements into code.
+(2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort.
+(3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
+The history of software development is characterized by a continuous pursuit of enhanced efficiency, superior quality, and greater automation capabilities.
+To better contextualize the significance of our proposed ”Token-Aware Coding Flow,” we first review and analyze the evolution of software development paradigms, particularly highlighting the recent transformations driven by Large Language Models (LLMs).
+As depicted in Figure 2, the evolution of software development models can be broadly categorized into the following three core stages[15][16][17]:
+II-A Manual Programming
+In the early phase of software development (Stage 1, Fig.
+2), the process relied almost entirely on manual efforts by programmers.
+Every step—from requirements to design, implementation, testing, and maintenance—depended on their expertise and time.
+While this approach maximized human creativity, it suffered from low efficiency, limited scalability, difficult knowledge transfer, and a high risk of human errors and inconsistencies.
+II-B Assisted Programming
+Advancements such as IDEs and code assistance tools ushered in an assisted programming era (Stage 2, Fig.
+2), where programmers leveraged technologies like auto-completion, syntax highlighting, version control, and basic code generators to boost productivity.
+More recently, LLM-powered tools (e.g., GPT-3.5, GitHub Copilot) have enabled smarter code suggestions, test generation, and documentation drafts.
+As a result, the programmer’s role shifted towards supervision and validation, marking the start of true human–AI collaboration.
+II-C Cognitive Programming
+Today, development is increasingly driven by advanced reasoning models (Stage 3, Fig.
+2).
+Powerful LLMs with Chain of Thought (CoT) capabilities (such as GPT-o1, DeepSeek-R1[1], GPT-4, DeepSeek Coder) can understand, decompose, and autonomously address complex requirements—planning, generating, and optimizing multi-step code, and even participating in architecture and debugging.
+Developers now focus on defining goals and making decisions, while models handle much of the cognitive and implementation workload.
+This paradigm greatly enhances efficiency and problem-solving, potentially transforming software engineering.
+However, as discussed in the Introduction, this reasoning-driven paradigm—especially its reliance on CoT—also leads to a sharp rise in token consumption.
+Optimizing token efficiency is thus essential to fully realize its benefits and ensure sustainable use in diverse development scenarios.
+This challenge motivates our proposal of the “Token-Aware Coding Flow” and the exploratory work in this study.
 Figure 3: Illustration of the LLM-driven code generation workflow in real-world development scenarios.
@@
 The process concludes with the production of executable code that satisfies the development goal.
-III Experimental Insights
+III Crafting the Future House
 III-A Research Questions
@@
 First, RQ1 primarily focuses on whether smelly code causes large models to consume more tokens during inference, which is crucial for evaluating the model’s ability to handle non-standard code.
-RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency.
-RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption.
-RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption.
+RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency[6][19].
+RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption[19].
+RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption[19].
 Finally, RQ5 explores whether other effective prompt engineering strategies, such as context awareness, role constraints, and cost-sensitive strategies, can further optimize reasoning efficiency beyond code refactoring.
@@
 III-B1 Observational Experiment
-Datasets: We used the Java dataset (Text-To-Code) from CodeXGLUE[14] for the experimental data, which includes 300 samples of smelly code and 300 samples of clean code.
-These datasets were used to analyze the behavior of the DeepSeek-R1 reasoning model when processing these codes, particularly the differences in token consumption and time consumption.
-Metrics: We recorded the total number of tokens consumed by the model for each code and the time taken for reasoning.
-The token consumption per unit time was calculated as the core metric to assess the efficiency of large models when handling different types of code.
-Additionally, the generated code was evaluated based on functionality correctness, readability, robustness, and other indicators.
-III-B2 Refactoring[22] and Optimization Experiment
-Datasets: In this experiment, we used the dataset based on smelly code and applied DeepSeek-R1 for code refactoring to generate refactored code (rf_code).
-The refactored code was used for chain processing to compare its reasoning efficiency with the original smelly code and clean code.
-Metrics: We recorded the number of tokens consumed in each processing step and calculated the Halstead complexity and cyclomatic complexity.
-By assessing the token consumption per unit of Halstead complexity and per line of code, we evaluated the efficiency improvements of the refactored code during model inference.
-Further, the accuracy of the generated code was assessed through functionality consistency scores (e.g., CodeBLEU) to ensure that the refactored code maintains functional consistency with the original code.
+Datasets: We used the CodeXGLUE Java (Text-To-Code) dataset, containing 300 smelly code samples and 300 clean code samples, to analyze differences in token and time consumption when DeepSeek-R1 processes different types of code.
+Metrics: We recorded the total number of tokens and inference time for each code sample.
+Token consumption rate was calculated to evaluate the model’s efficiency in handling different code types.
+The generated code was further assessed for functional correctness and readability.
+III-B2 Refactoring and Optimization Experiment
+Datasets: Smelly code samples were refactored using DeepSeek-R1 to generate refactored code, which was then compared with the original smelly code and clean code to assess inference efficiency.
+Metrics: We tracked token consumption at each processing step and calculated Halstead and cyclomatic complexity.
+Token consumption per unit of complexity and per line of code was analyzed.
+Functional consistency was ensured using metrics such as CodeBLEU.
 III-B3 Code Smell Type Comparison Experiment
-Datasets: We selected code samples from 10 different types of code smells in the smelly code dataset, with 50 samples from each type, totaling 500 samples.
-These samples were used to analyze the impact of different types of code smells on token consumption during model inference.
-Metrics: By comparing the token consumption during reasoning for different types of code smells, we calculated the token consumption per unit time for each smell type.
-The differences in the impacts of design, structural, and naming-related smells were further analyzed.
-All generated code was evaluated for functional consistency to ensure that the model-generated code met the expected functionality during reasoning.
+Datasets: We selected 10 types of code smells from the smelly code dataset, with 50 samples per type (500 samples in total), to analyze the impact of different smell types on token consumption.
+Metrics: Token consumption during inference was compared across different smell types, with additional analysis of design, structural, and naming smells.
+The generated code was also evaluated for functional consistency.
 III-B4 Explicit Smell Type Prompt Experiment
-Datasets: This experiment was conducted on the previously used smelly code dataset, primarily comparing the token consumption differences when explicitly indicating code smell types in the prompt versus not indicating them.
-Metrics: We recorded the number of tokens consumed in each experimental step and calculated token consumption per unit time.
-After generating the code, functionality consistency was assessed using metrics like CodeBLEU.
-Additionally, the effect of explicitly indicating code smell types was evaluated based on the model’s reasoning time, token consumption, and output quality.
+Datasets: Based on the smelly code dataset, we compared token consumption when the prompt explicitly indicated the code smell type versus when it did not.
+Metrics: Token and time consumption were recorded at each experimental step.
+Functional consistency of the output was evaluated using metrics such as CodeBLEU.
+The impact of explicitly indicating the code smell type on inference efficiency and code quality was also analyzed.
 III-B5 Prompt Optimization Strategy Experiment
-Datasets: In this experiment, we tested different prompt optimization strategies on the smelly code dataset[23], analyzing the effects of Context Awareness, Responsibility Tuning, and Cost-Sensitive strategies.
-Metrics: We recorded token consumption during the reasoning process for each strategy and calculated the Halstead complexity and cyclomatic complexity for each code.
-By comparing token consumption per unit complexity and per line of code, we assessed the impact of the three prompt optimization strategies on reasoning efficiency.
-All experiments were also evaluated for functionality consistency and code generation quality to ensure the generated code was accurate[16].
+Datasets: Various prompt optimization strategies—such as context enhancement, responsibility tuning, and cost-sensitive prompting—were tested on the smelly code dataset.
+Metrics: Token consumption and code complexity under each strategy were recorded.
+Token consumption per unit complexity and per line of code was compared.
+Each strategy was evaluated based on functional consistency and code generation quality.
 III-C Implementation
@@
 The flow demonstrates the interaction between different components, including code refactoring and requirement extraction, with the evaluation steps represented through similarity comparison.
-  • Data Processing [20] and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed with Tree‑Sitter to detect and classify instances of code smell.
-Leveraging Tree‑Sitter’s high‑performance syntax analysis, we annotated ten common smell categories, thereby constructing parallel “smelly” and “clean” code datasets.
-For each smelly sample, we invoked the DeepSeek‑R1 API to execute the reasoning task, recording both token consumption and inference latency.
-We then applied DeepSeek‑R1’s refactoring API to eliminate smells—producing rf_code—and compared the inference efficiency (tokens consumed and time taken) between original and refactored versions.
-  • Prompt Engineering and Optimization[25]:To further reduce token overhead, we evaluated three prompt‑level strategies: context awareness, role constraints, and cost sensitivity.
-Each strategy was realized by modifying the model’s input prompt, thereby influencing the reasoning trajectory and token usage.
-For every experimental condition, we issued the corresponding API calls to adjust generation parameters and logged token counts, inference durations, and the resulting code quality.
-  • Automation, Logging, and Analysis:All API interactions were fully automated via Python scripts, ensuring consistent and efficient execution.
-Each experiment’s outputs—including raw logs of token usage, timing, and generated code—were aggregated and processed by custom analysis scripts.
-These scripts perform data cleansing, statistical aggregation, and visualization, enabling systematic comparison across differing experimental configurations.
-IV Experimental Results
-IV-1 RQ1: Smelly Code vs. Clean Code
+  • Data Processing and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed using Tree‑Sitter to annotate ten common code smell categories.
+This enabled the construction of parallel “smelly” and “clean” datasets.
+For each sample, we recorded token consumption and inference latency during both code repair and refactoring tasks using the DeepSeek‑R1 API.
+  • Prompt Engineering and Optimization: To mitigate token overhead in CoT-based reasoning, we implemented and evaluated three prompt-level strategies: Context Awareness, Responsibility Tuning, and Cost Sensitive.
+Each strategy was applied by modifying the model’s input prompt, and for each experiment, we logged token usage, inference time, and code repair quality.
+  • Automation and Analysis: All API interactions were automated via Python scripts, ensuring reproducibility and efficiency.
+Outputs—including token counts, timing, and repaired code—were collected and analyzed using custom scripts for data aggregation, statistical analysis, and visualization, enabling comprehensive comparison of different strategies.
+IV Experimental Study: Triggering the nano surge
+Figure 5: Nano Surge on Token Consumption
+IV-A RQ1: Smelly Code vs. Clean Code
 Experimental Design.
@@
 Statistical evaluation reveals that smelly code consistently demands higher token expenditure(Fig.
-5)(Tab.
+6)(Tab.
 I).
@@
 These findings suggest that code smells not only harm maintainability but also increase the inference cost of large models, highlighting the importance of code quality for efficient model reasoning.
-IV-A RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
+IV-B RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
 Experimental Setup.
-To answer RQ2, we designed an experiment to investigate how code refactoring—specifically the elimination of code smells—affects token consumption during model inference.
-We first used the DeepSeek‑R1 inference model on both smelly_code.jsonl and clean_code.jsonl to establish baseline token‑usage data.
-Next, we applied automated refactoring to the smelly code samples, removing common smells (e.g., long methods, duplicated logic) to generate a refactored code set (rf_code).
-We then re‑ran inference on these refactored samples and compared their token consumption to the original smelly code.
-To ensure functional consistency, we employed CodeBLEU and other metrics to verify that the refactored code preserved the original behavior.
+To address RQ2, we examined how refactoring to remove code smells impacts token consumption during model inference.
+Using DeepSeek‑R1, we first established baseline token usage on both smelly_code.jsonl and clean_code.jsonl.
+We then applied automated refactoring to the smelly code (removing common smells like long methods and duplicated logic) to produce a refactored code set (rf_code), and measured token usage on these samples.
+Functional consistency was verified using CodeBLEU and related metrics.
 Analysis.
-During analysis, we recorded both the inference duration and token consumption for each trial and computed token consumption per unit time.
-Under the functional consistency constraint, refactored code is expected to be leaner, reducing unnecessary complexity and redundant patterns.
-Our results show that after refactoring, the DeepSeek‑R1 model performs inference more efficiently: despite maintaining equivalent functionality(Tab.
-III), the streamlined code requires significantly fewer tokens(Fig.
-6)(Tab.
+For each run, we tracked inference duration and token consumption, and normalized by time.
+Results show that refactored code leads to more efficient inference—maintaining functionality (Tab.
+III), but requiring substantially fewer tokens (Fig.
+7, Tab.
 II).
-In particular, refactoring complex smells—such as excessively long functions and duplicated segments—substantially reduced the number of reasoning steps and validations the model must perform.
+Refactoring complex smells, especially long functions and duplicate code, significantly reduced the number of reasoning steps and validations the model performed.
 Figure 7: Comparison of complexity-normalized token consumption before and after code refactoring.
@@
 These results suggest that systematic elimination of code smells is a critical strategy for reducing computational overhead in large‑model reasoning.
-IV-B RQ3: Impact of Different Types of Code Smells on Token Consumption
+IV-C RQ3: Impact of Different Types of Code Smells on Token Consumption
 Experimental Design / Setup
@@
 The first graph (Fig.
-7) and the table(Tab.
-IV))present the average growth rate of token consumption across various code smell types.
-It is evident that most of the code smells, such as binary_operator_in_name, complicated_boolean_expression, and complicated_regex_expression, lead to a significant increase in token consumption.
-These types of smells are often characterized by complex logical constructs that demand deeper reasoning from the model, resulting in more tokens being processed[26].
-Conversely, the smells func_name and too_long_parameter_list have a relatively lower impact on token growth.
-This suggests that these types of smells are simpler and less likely to introduce complexity in reasoning, causing fewer tokens to be consumed.
+8) and table (Tab.
+IV) present the average growth rate of token consumption for different code smell types.
+Most smells, such as binary operator in name, complicated boolean expression, and complicated regex expression, cause a notable increase in token usage, likely due to their complex logic requiring deeper model reasoning [22].
+In contrast, smells like func name and too long parameter list have a lower impact, indicating they are less complex and lead to only minor increases in token consumption.
 The second graph (Fig.
-8) categorizes code smells into four groups: Naming, Expression, Structure, and Design.
-The Expression category stands out with the highest average token consumption growth, followed by Structure.
-These categories typically involve complex logical constructs or convoluted code structures, which require more tokens for the model to parse and reason about.
-In contrast, the Naming and Design categories exhibit lower token consumption growth.
-Issues like poor naming conventions and design flaws might affect code readability and maintainability, but they do not significantly complicate the reasoning process, leading to lower token usage by the model.
+9) groups code smells into Naming, Expression, Structure, and Design.
+Expression-related smells show the highest token consumption growth, followed by Structure.
+These categories often involve intricate logic or convoluted code structures, demanding more tokens for the model to process.
+By comparison, Naming and Design smells have a smaller effect on token growth, as they are less likely to complicate the reasoning process.
 Figure 8: Average token growth rate per code smell type.
@@
 These findings suggest that prioritizing the removal of design and structural smells can maximize inference efficiency and substantially reduce computational overhead.
-IV-C RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[27]
+IV-D RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[23]
 Experimental Design / Setup
@@
 Results (Fig.
-9)(Tab.
+10)(Tab.
 VI) indicate that adding explicit smell annotations to the prompt yields a more efficient reasoning process, with a marked reduction in token usage.
@@
 In contrast, without smell hints, the model must engage in multiple verification cycles to detect and interpret hidden code complexities, thereby inflating token consumption.
-Figure 9: Comparison of total token consumption for smelly code with and without explicit prompt tips
 TABLE VI: Descriptive Statistics of Total Token Consumption for Smelly Code (With vs. Without Prompt Tips)
@@
 Max 16440.00 15045.00
+Figure 10: Comparison of total token consumption for smelly code with and without explicit prompt tips
 TABLE VII: Comparison of CodeBLEU and Similarity Metrics with and without Tips
@@
 This strategy thus significantly enhances inference efficiency and mitigates resource usage in large‑model code reasoning.
-IV-D RQ5: Prompt Engineering Strategies beyond Refactoring
+IV-E RQ5: Prompt Engineering Strategies beyond Refactoring
 TABLE VIII: Mean Values of Key Metrics under Different Strategies
@@
 To address RQ5, we designed a series of experiments(Fig.
-10) aimed at evaluating three prompt engineering strategies—Context Awareness, Responsibility Tuning, and Cost Sensitivity—and their effect on reducing token consumption during the model’s reasoning over smelly code.
+5) aimed at evaluating three prompt engineering strategies—Context Awareness, Responsibility Tuning, and Cost Sensitivity—and their effect on reducing token consumption during the model’s reasoning over smelly code.
 Each strategy modifies the input prompt to guide the model’s cognitive process, driving a more efficient and focused inference, while minimizing unnecessary token expenditure.
@@
 By constraining the token size, we controlled the model’s computational cost, particularly when generating larger code segments, effectively reducing token expenditure without sacrificing the integrity of the generation.
-All experiments were performed on the same dataset of 300 smelly code snippets (smelly_code.jsonl).
+All experiments were performed on the same dataset of 300 smelly code snippets.
 For each condition, we invoked DeepSeek‑R1, recording the total token consumption and calculating token consumption per unit time.
@@
 However, overly stringent constraints sometimes led to the truncation of essential code elements, resulting in reduced functional correctness and lower accuracy in the output.
-Figure 10: Nano Surge on Token Consumption
-Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning emerged as the most effective strategies for reducing token consumption beyond refactoring(Tab.
+Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning proved most effective for reducing token consumption beyond refactoring (Tab.
 IX).
-Context Awareness decreased token usage by 15–20%, especially on lengthy or complex code.
-Responsibility Tuning achieved a 10–15% reduction by framing the model’s reasoning role.
-Cost Sensitivity, while reducing tokens by 20–30%, risked incomplete code generation under strict limits.
-These results suggest that combining context-rich prompts with precise role framing offers substantial efficiency gains without compromising functionality.
-Furthermore, the optimizations applied through these strategies also led to improvements in code quality metrics.
-Specifically, the codeBLEU score increased by 38%, and function similarity was enhanced by 4%, compared to the baseline performance.
-These improvements indicate that not only were token consumption and reasoning efficiency optimized, but the overall quality of the generated code also saw significant gains, further validating the efficacy of these prompt engineering strategies.
-V Related Work
+Context Awareness reduced token usage by 15–20%, especially for complex code, while Responsibility Tuning achieved a 10–15% reduction by specifying the model’s reasoning role.
+Cost Sensitivity reduced tokens by up to 30% but sometimes caused incomplete outputs.
+Overall, combining context-rich prompts with clear role framing offered notable efficiency gains without sacrificing functionality.
+These optimizations also improved code quality: codeBLEU scores increased by 38%, and function similarity improved by 4% compared to the baseline, demonstrating that prompt engineering can simultaneously enhance efficiency and output quality.
+V Discussion
+This study has conducted preliminary observations and analyses of the token consumption behavior exhibited by Large Language Models (LLMs) in Chain of Thought (CoT) scenarios[4].
+The findings reveal complex interactions between code characteristics and token overhead, alongside several possibilities for optimization through external strategies[24].
+The specific behavior of LLMs in software engineering tasks, particularly their resource consumption patterns (such as token usage), often demonstrates a high degree of contextual dependency and certain ”black box” characteristics, rendering precise prediction solely through theoretical models challenging.
+In this context, the long-standing empirical tradition within software engineering—which emphasizes understanding and mastering complex systems through iteration, observation, experimentation, and data-driven analysis—offers a potent perspective and methodology for addressing such challenges[25].
+When confronting the specific issue of token consumption by LLMs in code processing, an approach grounded in empiricism is particularly crucial[26].
+Empirically observing the actual impact of various factors, such as different code features and prompting strategies, on token consumption, and quantitatively evaluating the effectiveness of intervention measures (like the code refactoring and prompt optimization explored in this study), serves as an effective pathway to deeply understand and manage this novel form of ”resource consumption.” The analytical framework of the present study represents a preliminary endeavor guided by this philosophy.
+Looking ahead, the role of LLMs in software development and associated areas of focus are expected to continually evolve.
+On one hand, as the capabilities of LLMs iteratively improve, the conformity of the code they generate, or their ability to understand and intelligently refactor existing code[12], is poised for significant enhancement.
+This may gradually reduce the future need for manual attention to, and remediation of, traditional ”code smells,” potentially shifting the focus of software quality assurance to new dimensions.
+On the other hand, even if LLMs can efficiently process or generate ”smell-free,” high-quality code, the reasoning and generation processes essential for fulfilling complex requirements will inherently continue to incur token costs[27].
+Therefore, regardless of how the issue of ”smells” at the code level evolves, attention to token consumption will remain a core concern in the field of LLM-assisted software development.
+Optimizing the token efficiency of the generation process and striking the optimal balance between output quality, functional completeness, and resource expenditure will continue to be key challenges and top priorities for future research aimed at the sustainable application of this technology[28].
+VI Related Work
 Over recent years, with the widespread application of large language models (LLMs) in natural language processing (NLP)[4], automated code generation and optimization have also made significant strides[29].
 Research in this area can be broadly categorized into three directions:
-V-A Code Smell Detection and Refactoring
+VI-A Code Smell Detection and Refactoring
 Code smells are syntactic or structural code features that do not immediately impair program functionality but degrade readability, maintainability, and extensibility[30][12][13].
@@
 Nevertheless, efficiently handling complex smells remains challenging, particularly in deep learning–based inference where token consumption and computational overhead can escalate.
-V-B LLM‑based Code Generation and Optimization
+VI-B LLM‑based Code Generation and Optimization
 The advent of pre‑trained language models has spurred extensive research into LLM‑driven code generation.
@@
 To mitigate this, researchers have explored integrating automated refactoring with LLM training to jointly optimize code quality and inference efficiency.
-V-C Integration of Code Quality and Inference Efficiency
+VI-C Integration of Code Quality and Inference Efficiency
 Recently, there has been growing interest in coupling code quality improvements with inference efficiency in pre‑trained models.
@@
 Building upon this work, our study investigates how code smell removal and prompt engineering jointly reduce token overhead, particularly on complex code fragments.
-VI Conclusion and Future Work
+VII Conclusion and Future Work
 This study investigated strategies to optimize the reasoning pipeline of large language models (LLMs) in order to reduce token consumption when processing code with smells[14][32].